Instrument Log

 

Disdrometer/ORG/WRG      C1     Lamont, OK

 

Date

Technician

Comments

7/15/2016

MS

The ORG Instrument was not running during PM rounds this morning.  Power cycled the data logger but to no avail.  The data logger has a heartbeat but has not collected files since 7-13-2016.  Still trouble shooting instrument.

7/14/2016

MS

The ORG Instrument was not running even after replacing the battery in the data logger on 7-13-2016.  At 2100 GMT replaced big battery in the Power Supply due to very low voltage,  program started back up.   RTS at 2115 GMT.

7/13/16

MS

The Disdrometer Adjusted Server time was 20 seconds behind the Station time.  Set the Adjusted Server Date/Time to match the Station Date/Time to July 13, 2016  at 13:24:00.  RTS 1328 GMT.

07/13/2016

MS

The ORG program had stopped running at 1323 GMT during the time of resetting the Adjusted Server time clock on the Disdrometer.  At 2000 GMT replaced the battery in the data logger and started program back up.  RTS 2015.

6/6/2016

MS

The ORG had stopped working on 6/3/2016 at 1900 GMT.  At 1306 GMT tried to shut down the instrument/program and bring back up.  Didn’t work so went to the power box and unplugged instrument at 1315 GMT and then waited for 15 seconds and plugged instrument back in.  RTS at 1330 GMT.

6/3/16

KT

Work order requested that the WBRG be removed from service and placed down into storage until further notice. At 1656 GMT the program and collections were turned off and the instrument was removed from service. Model # AEPG600, SN# 00073, WD# 59485. RTS at 1900 GMT.

4/25/16

KT

DQPR 5240: Disdrometer is not reporting rainfall accurately, recording data only for the first few minutes of every hour.  Mentor requested replacing the operating disdrometer with the one that was returned from Darwin.  At 1614 GMT the program was stopped and the Disdrometer was removed and replaced with the one from Darwin. At 1830 GMT powered system on.

*Old - 2383405  *New - 2393405

1/27/2016

MS

Instrument is down as of 1600 GMT, 1/26/2016 for a scheduled Virtual Computer Host upgrade.    Data will be lost until upgrade is complete.

1/6/16

JEM

The time on the WBRG  datalogger is 46:08:59 slow and needs to be set.  Current time was 01/06/2016  21:05:20 GMT.  The datalogger time was showing 01/04/2016  19:56:21 GMT.  Set the clock to the proper time at 01/06/16  21:06:00 GMT.  All data that has been plotted since 12/28/2015 at approximately 1345 GMT until the time was properly set is incorrect.  RTS @2115 GMT.

1/6/16

JEM

The time on the ORG datalogger is 46:08:59 slow and needs to be set.  Current time was 01/06/2016  21:05:20 GMT.  The datalogger time was showing 01/04/2016  19:56:21 GMT.  Set the clock to the proper time at 01/06/16  21:06:00 GMT.  All data that has been plotted since 12/28/2015 at approximately 1345 GMT until the time was properly set is incorrect.  RTS @2115 GMT.

12/15/15

MS

Computer needed to be restarted following scheduled software and security patch updates.  At 1404 GMT computers were updated and rebooted by the IT department, some data might be lost. 

11/5/15

JEM

Sensor #2 continues to go offline causing the frequency and weight for the sensor to flat line at zero.  R/R *sensor #2.  After replacing the sensor, edited the WBRG configuration to reflect the change @1520.  New calibration information is T=24.91, C1=2.55804, C2=0.01096 and TC=14.30.  After the replacement, the readings were normal and in line with the other 2 sensors.   RTS @1530 GMT.

*Old – 666,  New – 529

9/16/15

 

The computer needs to be rebooted following scheduled software and security patch updates.  Stopped program at 1513 GMT and rebooted computer. Restarted and RTS at 1530 GMT.

8/24/15

JEM

Looking at the data plots for the WBRG I noticed that sensor #2 has been dropping out. Troubleshooting found a poor connection where the sensor plugs onto the electronics board.  Secured this connection and re-zeroed the rain gauge @1745 GMT.  RTS @1800 GMT.

8/24/15

JEM

DQPR #4849 indicates the ORG is reporting large amount of precipitation with no rain actually occurring and the IR Carrier Level is also fluctuating.  Troubleshooting found the battery and power supply to be very hot.  When the battery was removed from the power supply it only measured 11.5V.  Suspect the battery is beginning to fail.  R/R the 12V battery and the power supply @1650 GMT.  RTS @1655 GMT.

8/27/15

CAM

No data had been collected since 0105 GMT.  Logged in remotely and found Disdrodata program was running, but LoggerNet was connected, possibly preventing SDS to connect for collections.  Closed LoggerNet at 1339 GMT and RTS.

8/4/15

JEM

DQPR #4830 for WBRG indicates "All precip/weight variables flatline starting 7/29. The total_mm variable drops out to 0 as well."  Connected to the rain gauge with a laptop and observed the total mm at exactly 0. Pressed on the bucket slightly and the total mm remained at zero. Sent an 'rZ' command to the rain gauge to zero it again. Pressed on the bucket slightly and the total mm climbed to ~80, then slowly fell back to ~0. After the bucket stabilized, issued the 'rZ' command to zero the gauge again. Gauge appears to be operating normally now. RTS @2035 GMT.

7/29/15

KT

The weighing rain bucket needed to be emptied. At 1740 GMT engaged port flag 1 and emptied the weighing rain gauge bucket. At 1755 added new fluids to the bucket and installed. At 1803 disengaged port flag 1. The reading was at 0. RTS at 1805 GMT.

7/8/15

JEM

DQPR #4780 indicates the Disdrometer is not recording any rain data.  Tapped on the cone and pressed the test button.  Neither one showed up on the disdrodata program.  Power cycled the processor and then tapped on the cone and pressed the test button again.  Still, neither showed up on the disdrodata program.  Connected a loopback to the MOXA in the disdrometer control box.  Was unable to loopback any data using Putty.  Tried to ping the MOXA and got no response.  Inspection of the MOXA revealed that it was not linking with the network hub.  R/R the MOXA and set the Disdrodata software to COM3 before starting data collection.  After the program started, tapped on the cone and pressed the test button.  Both were recorded in the data at 1625 GMT.  RTS @1630 GMT.

7/6/15

JEM

Due to a power outage overnight, the disdrometer computer restarted and the software did not restart.  Added Disdrodata v1.7 to the startup folder.  Opened Disdrodata and checked the box for automatically starting data collection.  Started data collection @1317 GMT.  The Disdrometer returned to normal operation.  RTS @1325 GMT.

6/22/15

CAM

No data had been collected from the disdrometer since 0005 GMT on 6/20.  Checked status and found Disdrodata program was not running.  Started Disdrodata program and RTS at 1511 GMT.

6/17/2015

MS

Disdrodata program had stopped running for the Disdrometer at 1000 GMT.  Restarted Disdrodata program for the Disdometer at 1301 GMT. RTS at 1316 GMT.

6/10/2015

MS

Due to a scheduled power outage the Disdrometer program was shut down at 1353 GMT.  Restarted Disdrometer program at 1601 GMT.  RTS at 1616 GMT.

6/4/2015

MS

Shutdown Disdrometer programs at 1322 GMT due to scheduled power outage.  Restarted Disdrometer programs at 1848 GMT.  RTS 1903 GMT.

5/12/15

JEM

While checking the data to see if recent rain events were properly recorded, sensor #1 was observed flatlined at zero for both weight and frequency.    Logged into the WBRG computer remotely and both readings were NaN and the sensor 1 temperature was -80C.  Physically inspected the instrument and found that sensor #1 has completely separated into two pieces.  R/R *sensor #1 with a spare sensor.  Updated the calibration file with the proper coefficients for the new sensor.  Performed both zero and span calibration.  Re filled the bucket with the proper amount of water and oil and zeroed the bucket again.  RTS @1810 GMT.  *Old – 529  *New – 619

5/7/15

JEM

After having a conference call with the mentor and Topper from Belfort, the calibration process needs to be repeated and configuration files needs to be changed.  Changed the expected mm of rain in the program to the correct number provided by Belfort.  Performed both zero and span calibrations.  Filled the bucket with the proper amount of water and oil.  Zeroed the bucket again.  Dumped 500mL of water into the gauge and verified the reading at approximately 15mm as expected.  Zeroed the bucket again and returned to service.  RTS @1530 GMT.

5/6/15

JEM

Received the electronics and sensors to be installed in the WBRG.  Also received a span calibration kit to perform a span calibration and zero calibration on the gauge after the sensors are installed.  Installed the sensors and *electronics and powered up the system.  Performed a span calibration the best we could with the provided directions.  Filled the gauge with the proper amount of water and oil.  Zeroed the gauge and returned it to service @2045 GMT.  *New - 73

 

4/21/15

CAM

The computer needs to be rebooted following scheduled software and security patch updates.  Stopped program at 1827 GMT and rebooted computer. Restarted and RTS at 1830 GMT.

4/1/15

JEM

The ORG has not collected since 3/30/15 @1501 GMT.  Connected to the datalogger and only the data in the status table is updating.  Tried to format the datalogger and reload the program, but the data tables still would not update.  Loaded the latest OS onto the datalogger and reloaded the program.  Data tables would still not update.  R/R the *datalogger.  Was unable to load the program to the new datalogger because the OS on it was too old (OS06).  Updated the datalogger to OS28.  Loaded the program @1427 GMT.  Data tables are updating as expected and data appear reasonable.  RTS @1430 GMT.

*Old – 30068  *New – 2943

4/1/15

JEM

The WBRG *electronics and 3 strain sensors need to be returned to the vendor (Belfort) for electronics firmware upgrade and replacement of sensor #1 using RMA #B1511.  Removed the electronics and sensors from the WBRG @1454 GMT.  The WBRG will remain out of service until these items return from Belfort.

*Electronics - 73

Sensor 1 - 368

Sensor 2 - 373

Sensor 3 - 380

3/12/15

MS

Program was not connected at 1311 GMT.  Reconnected program at 1326 GMT.

2/10/15

JEM

SDS has been unable to collect from the ORG and WBRG for 9 hours.  Troubleshooting found the central cluster UPS had failed.  This caused the MOXA fiber hub to lose power.  R/R the UPS batteries.  Powered the UPS back on and communications links were restored at 1600 GMT.  RTS @1600 GMT.

2/9/15

JEM

The SDS has been unable to collect data from the ORG and WBRG since 0200 GMT on 2/8/15.  Troubleshooting determined that the data logger had failed.  R/R the *datalogger.  Loaded the logger program and verified operation at 1554 GMT.  RTS @1600 GMT. 

*Old – 15493   *New – 30068

12/16/14

CAM

No data had been collected from Disdrometer since 1805 GMT.  No data collected from ORG since 1601 GMT and WBRG since 1620 GMT.  As per BCR-02040, all SGP Windows computers were updated and rebooted today.  This required a manual restart of Disdrodota software.  Restarted Disdrodata program and Loggernet and RTS at 2159 GMT. 

10/30/14

CAM

No data had been collected from the disdrometer since 2305 GMT on 10/29.  Logged in remotely and found Disdrodata program had stopped updating and an error message was displayed “Error 16408 occurred at Serial Port Read.  Possible Reasons: Windows GetCommError.  The hex value of this code is x4018”.  Cleared error and Disdrodata program then scanned through all data from the time of shutdown up to current time.  Closed program and restarted.  RTS at 1316 GMT.

10/20/14

KT

SDS reported no files from the instrument. At 1238 GMT found Error 16408 occurred at serial port read. Cleared the error and restarted the program. RTS at 1253 GMT.

10/6/14

CAM

No data had been collected from the disdrometer since 2205 GMT on 10/5.  Logged in remotely and found Disdrodata program had stopped updating and an error message was displayed “Error 16408 occurred at Serial Port Read.  Possible Reasons: Windows GetCommError.  The hex value of this code is x4018”.  Cleared error and Disdrodata program then scanned through all data from the time of shutdown up to current time.  Closed program and restarted.  RTS at 1321 GMT.

10/1/14

CAM

Disdrodata program stopped updating at 1040 GMT.  Error message displayed:  Error 16408 occurred at Serial Port Read.  Possible reasons: Windows GetCommError, code x4018.  Cleared error and restarted program at 1318 GMT.

9/24/14

CAM

Disdrodata program stopped updating at 1512 GMT.  Error message displayed:  Error 16408 occurred at Serial Port Read.  Possible reasons: Windows GetCommError, code x4018.  Cleared error and restarted program at 1538 GMT.

9/24/14

CAM

Daily checks found Disdrodata had stopped updating at 0759 GMT.  Error message displayed:  Error 16408 occurred at Serial Port Read.  Possible reasons: Windows GetCommError, code x4018.  Cleared error and restarted program at 1230 GMT.

9/23/14

CAM

No data had been collected from the disdrometer since 2305 GMT on 9/22.  Logged in remotely and found Disdrodata program had stopped updating and an error message was displayed “Error 16408 occurred at Serial Port Read.  Possible Reasons: Windows GetCommError.  The hex value of this code is x4018”.  Cleared error and Disdrodata program then scanned through all data from the time of shutdown up to current time.  Closed program and restarted.  RTS at 1230 GMT.

9/8/14

CAM

No data had been collected from the disdrometer since 2205 GMT on 9/5.  Logged in remotely and found Disdrodata program had stopped updating and an error message was displayed “Error 16408 occurred at Serial Port Read.  Possible Reasons: Windows GetCommError.  The hex value of this code is x4018”.  Cleared error and Disdrodata program then scanned through all data from the time of shutdown up to current time.  Closed program and restarted.  RTS at 1231 GMT.

9/2/14

CAM

No data had been collected from the disdrometer since 8/29.  Logged in remotely and found Disdrodata program had stopped updating and an error message was displayed “Error 16408 occurred at Serial Port Read.  Possible Reasons: Windows GetCommError.  The hex value of this code is x4018”.  Cleared error and Disdrodata program then scanned through all data from the time of shutdown up to current time.  Although data was current, it did not appear reasonable considering it has been raining for several hours.  Closed and restarted program at 1244 GMT and data then seemed reasonable.

8/18/14

CAM

No data had been collected from the disdrometer since 0905 GMT.  Logged in remotely and found Disdrodata program had stopped updating and an error message was displayed “Error 16408 occurred at Serial Port Read.  Possible Reasons: Windows GetCommError.  The hex value of this code is x4018”.  Cleared error and restarted Disdrodata program at 1238 GMT.  The following collection was successful at 1405 GMT.

8/14/14

CAM

No data had been collected from the disdrometer since 0105 GMT.  Logged in remotely and found Disdrodata program was running.  LoggerNet was open and connected to ORG/WBRG datalogger with error message shown: Access violation at address 004892F4 ‘ connectscreen.exe’.  Read of address 00000004.  Disconnected LoggerNet and closed Disdrodata program and restarted at 1215 GMT.  The following collection was successful at 1305 GMT.

6/10/14

CAM

No data had been collected from the disdrometer since 2105 GMT on 6/9.  Logged in remotely and found Disdrodata program was running but scrolling rapidly through the data from start of today up to current time and then stopped and resumed normal operation.  LoggerNet was open and connected to ORG/WBRG datalogger.  Disconnected LoggerNet and restarted Disdrodata program.  RTS at 1214 GMT.

6/9/14

CAM

No data had been collected from the disdrometer since 0305 GMT on 6/7.  Logged in remotely and found Disdrodata program was running but scrolling rapidly through the data from start of today up to current time and then stopped and resumed normal operation.  Data files were current in dis_dir/out directory. LoggerNet was open and connected to ORG/WBRG datalogger.  Forced a collection from datalogger and then disconnected LoggerNet at 1223 GMT.  The following Disdrometer collection still failed at 1305 GMT.  Logged in remotely and closed Disdrodata program and restarted at 1337 GMT.  The following collection was successful at 1405 GMT.

5/28/14

JEM

The ORG has returned and needs to be reinstalled.  Installed and connected the *ORG sensor @2115 GMT.  Verified operation and RTS @2120 GMT.   *New - 05090146

5/19/14

CAM

No data had been collected from the disdrometer since 1305 GMT on 5/17.  Logged in remotely and found Disdrodata program was running but creating one large data file that started back on 5/17.  There were also several 1kb files created on 5/16 that contained only one sample each.  Stopped Disdrodata program and restarted at 1230 GMT.  Data collections returned successful at 1405 GMT.

5/1/14

CAM

No data had been collected from the disdrometer since 0905 GMT on 4/30.  Found error message displayed: Error 16400 Occurred At Serial Port Read.  Cleared error message and Disdrodata program started scrolling through all missing data until display was current.  Program then resumed normal operation.  RTS at 1208 GMT.

4/21/14

CAM

No data had been collected from the disdrometer for 12 hours.  Found error message displayed: There is a problem with Com3 (error code 37).  Cleared error message and Disdrodata program was set to resume at 1300 GMT.   Confirmed program started as it should at that time.  RTS status O.K.

3/10/14

CAM

No data had been collected from WBRG since 2021 GMT on 3/7.  LoggerNet was not running.  Opened LoggerNet, connected to datalogger and collected data. All missing data was transferred.  RTS at 1250 GMT.

11/26/13

JEM

DQPR #3898 indicates Sensor status for WBRG transducer 1 indicates a potential problem. Sensor is very noisy.  Set port 1 via Loggernet at 1529 GMT. Removed the WBRG cover and checked/reseated all transducer connections. The plug for transducer #1 had a large amount of moisture/mineral oil on it. Cleaned the moisture/oil off of the connector and blew the connector with canned air. No corrosion or rust was observed. LED status indicator for sensor #1 is also not lit. Transducers #2 and #3 are producing an audible tone, but transducer #1 is not. The WBRG was nearly full so we also emptied and re-zeroed it. This made the 3 sensors read more closely to each other, but when monitoring the WBRG with Loggernet, sensor #1 is still significantly more noisy that the other 2 sensors. The problem does not appear as severe now, but it may be proportional to the amount of weight in the bucket. Turned port 1 off @1625 GMT and RTS. Suspect that transducer #1 has failed.  RTS @1625 GMT.

11/15/13

JEM

The impact distrometer has been returned from the vendor and needs to be installed.  Installed the *distrometer and started the software collection program at 2108 GMT.  Installed the new version of distrodata that came with the instrument, but did not use it at this time.  Instrument mentor should look at the new software and make sure it will not alter the data stream in any way.  RTS @2110 GMT.  *New - 2383405

10/31/13

JEM

The ORG needs to be pulled and returned to the vendor for repair.  Removed the ORG @1535 GMT.  Notified the SDS team so they could disable ORG collections.   

9/4/13

CAM

The ORG has returned from being repaired.  Installed *ORG at 1815 GMT.  Output was 200mV.  Jumper between datalogger 1L and ground was missing.  Installed jumper and RTS at 1900 GMT.  Output was then 75mV and carrier voltage was 4.6V.

*New – Model ORG-815DA, SN 05090146

8/14/13

CAM

Stopped Disdrodata program at 1245 GMT and removed *instrument, processor, power supply and data cable from service to be returned to vendor for repair.

*Old - 2383405

8/13/13

CAM

Powered instrument off at 2110 GMT.  Will remove tomorrow to send to vendor for repair.  (Note: Disdrodata program was left running, so data may have continued to be collected overnight)

8/12/13

CAM

After not recording rainfall since 7/21, the disdrometer started working during rain event on 8/9 and also recorded rainfall during the 2 rain events since then.  Set flag #1 at 1813 GMT.  Tapped the cone several times and also pressed the test button to see if drops would be recorded.  Removed the cone to see if plunger movement could be detected.  Still unable to visibly or physically detect plunger movement but drops were recorded for all tests performed.  Removed flag #1 at 1819 GMT.  Notified mentor of findings.  Mentor suggests we still send the instrument to the vendor for repair and calibration.  Will remove instrument when RMA number is received from vendor.  *Note: data quality team later reported that the disdrometer recorded nearly twice as much rainfall as other instruments for each rain event after it started working on 8/9, so instrument was not working properly.

8/5/13

CAM

DQPR 3822 indicates the disdrometer has not reported rainfall since 7/21/13.  Pressing the test button results in LED #4 lighting on the processor as it should, but no drops are indicated in Disdrodata program.  Likewise, no drops are indicated when the cone is tapped.  Removed the cone from the plunger (but didn’t remove the rubber seal below it) and tried pressing the plunger to check for movement.  No visible movement detected.  Notified mentor, who suggested returning the instrument to vendor for repair.  Will create a PR and remove instrument once the PO and RMA number are assigned. 

7/17/13

CAM

DQPR 3733 indicates the ORG has not been comparing well with other rain sensors dating back to February, and is now reporting constant rainfall starting on 6/24.  Measured ORG output as 220 mV, which results in the constant 1.3 mm/hr of rainfall that is being reported.  Disconnected the jumper between the signal common and ground as well as the carrier common and ground and the signal output then measured -12V.  Disconnected heater +12V wire from datalogger and signal level then varied between -90 and -300 mV.  Measured impedance between heater +12 volt wire (white) and signal grounds as 30 ohms, which is normal according to OSI engineer.  However, the heater ground and signal grounds are not tied together internally as they should be.  Tried a new cable and verified problem is internal to sensor and not the cable.  Vendor suggested returning the sensor for repair and calibration.  Removed *ORG at 1930 GMT to return to vendor.

*Old – Model ORG-815DA, SN 05090146

6/13/13

CAM/PAD

WBRG bucket needs to be emptied and Tmm re-zeroed.  Turned port 1 on at 1817 GMT.  Emptied and cleaned bucket.  Added 1L of mineral oil to bucket and installed in cradle.  Reset Tmm to zero and installed plastic shield surrounding bucket.  Port 1 had already reset itself upon reconnecting to dataloger.  RTS at 1900 GMT.

5/28/13

CAM

DQPR 3733 indicates the ORG has not been comparing well with other instruments on site, usually reading much higher.  Checked all wiring, especially the grounds. No problems found. Measured ORG output as 50mV, which is acceptable during non-precip. Measured carrier voltage as 4.8V, which is in the acceptable range. Reseated all connections at 1620 GMT.  Will see how ORG compares during next rain event.

5/28/13

CAM

Data collections for ORG, WBRG (RAIN) and disdrometer started failing at 0901 GMT today.  The drive on the virtual computer had completely filled up over the weekend with 0% of 8 GB available.  T/S found a data file, LAN_2_CR1000_DataSample, in C:\Campbellsci\LoggerNet\ that was 1.79 GB in size, current, and growing.  Went into LoggerNet setup and disabled the DataSample table entry under the Data Files tab.  Copied the large data file to another computer and deleted the original file.  This freed up 1. 79 GM on the virtual drive.  Restarted programs and RTS at 1544 GMT.  The following data collections were successful.

5/22/13

CAM

No data had been collected from the ORG or WBRG for 4 hours.  Logged into computer remotely and found Disdrodata program also displayed an error writing to current file.  Attempted to restart Disdrodata but there was insufficient disk space to start LabView.  Checked disk space and found there was only 332 KB free on the 8 GB drive.  Checked all data directories to see if there was old data that was not being deleted, but none found.  Did find 2 identical directories containing the instrument handbook, so deleted one of them, which freed up 22 MB of disk space.  This is still extremely low, but should allow the programs to run until IT department can choose what additional content to delete from hard drive.  Restarted Disdrodata and RTS at 1930 GMT.  Notified IT department that more space will need to be freed up. 

2/26/13

CAM

The site lost power at approximately 0350 GMT on 2/26 due to a severe winter storm.  Powered central cluster UPS off at 1730 GMT.  Power was restored at 1920 GMT.  Powered UPS on at 1945 GMT.  Instrument RTS status O.K and data collections resumed once SDS operations were restored.

11/20/12

CAM

Local electrical service was shut off at 1410 GMT so that Kay Electric could upgrade its power lines from 7.2 KV to 14.4 KV in Grant County.  Disdrometer was gracefully shut down at 1348 GMT in anticipation of power outage.  Power was restored at 2115 GMT after electrical upgrades were complete.  Restarted disdrometer and RTS at 2300 GMT.

11/13/12

CAM

Site power was partially shut down at 1400 GMT in order to replace the 7 transformers that service the buildings and all instruments south of the main compound.  Power was restored at approximately 1930 GMT after all transformers were replaced.  Disdrometer computer displayed error with comm port 3 (error code 37).  Cleared error and restarted Disdrodata and Loggernet programs.  RTS at 2122 GMT.

10/11/12

CAM

The disdrometer has been returned from the vendor after being repaired and needs to be installed.  Installed *disdrometer and powered on at 2135 GMT.  Depressed test button and observed LED #4 light up as expected.  Notified mentor that instrument was installed.

*New – 2383405  Model RD-80

7/20/12

JEM

The impact disdrometer is not recording precipitation and needs to be returned to the vendor for repair.  Powered the *disdrometer off @1723 GMT.  Removed the processor, *sensor and cable and took it to shipping to be returned to the vendor.  *Old - S/N - 2383405, model RD80

6/20/12

PAD

DQPR 3443 on 6/18/12 states "The instrument is not recording precipitation while other instruments are".  Mentor has determined 6/12/12 was date instrument failed.   Previous troubleshooting found no movement when cone was tapped or pressed.  Mentor requests cone and rubber seal beneath cone removed and plunger inspected.  Set flag on port 1 at 1549 GMT.  Removed cone and seal and inspected what was visible of plunger.  Could see no movement where expected when plunger was pressed.  Took pictures of instrument with cone and seal removed and sent to mentor.  Troubleshooting with mentor at this time.  Removed flag at 1616 GMT.  RTS at 1619 GMT. 

6/19/12

PAD

DQPR 3443 on 6/18/12 states "The instrument is not recording precipitation while other instruments are".  Mentor requests test button pressed on processor and plunger tapped to see if drops register on disdrodata.  Set flag on port 1 at 2022 GMT.  Pressed processor test button and number 4 light on processor lit up while button was pressed.  Tapped on plunger.  Nothing was shown from either test on disdrodata program.  Shut down/ restarted processor and set program to start measurements at 2100 GMT.  Repeated tests at 2102 GMT.  No change in status.  Troubleshooting with mentor at this time.  RTS at 2105 GMT.

4/17/12

PAD/KT

WBRG bucket needs to be emptied and Tmm re-zeroed.  At 1811 GMT, set Port 1 on datalogger.  Skimmed mineral oil off top of bucket to reuse.  Emptied and rinsed bucket.  Added approximately 2L of water and mineral to bucket and installed in cradle.  Reset Tmm to zero and installed plastic shield surrounding bucket.  RTS at 1851 GMT.

3/28/12

PAD

Observed disdrometer computer clock is one hour fast.  At 1252 GMT, changed time zone from GMT to GMT Casablanca and unchecked "adjust for daylight savings time".  Computer clock is now correct.  RTS at 1307 GMT.

11/9/11

PAD

ORG is sporadically showing false rain fall.  Mentor requests troubleshooting for possible ground loops causing false readings.  At 2002 GMT, display shows ORG_volt measures 44 mV.  Disconnected ORG chasis ground wire and ground wire connecting chassis ground and datalogger ground rods.  ORG_volt measured 79mV.  ORG_volt measures .79 mV with ORG chassis ground connected and wire between ground rods disconnected.  Disconnected datalogger ground wire and ground rod connecting wire, ORG_volt measures .79 mV.  With datalogger connected to ground ORG_volt measures .79 mV.  Reconnected all grounds and wire connecting ground rods and reading returned to 44 mV.  RTS at 2017 GMT.

10/18/11

PAD

On 10/11/11, observed WBRG values are not shown on Numeric 2 display.  Connected laptop directly to instrument but no data is scrolling.  Found loose connection at datalogger wiring panel on serial cable coming from instrument.  Serial cable was disconnected from wiring panel on 10/6/11 to allow password to be entered and make parameter change mentor had requested.  Serial wire connector was not seated correctly.  Reseated serial wire and data started scrolling on laptop and is now shown on Numeric 2 display.  RTS at 1525 GMT.

10/6/11

PAD

DQPR 2732 states for WRG "Please change the setting in the configuration file for Averaging Interval from the manufacturer's default of 10 seconds to 60 seconds".  Connected with WRG and made change to parameter file as requested.  RTS at 1902 GMT.

10/3/11

PAD

ORG instrument has been returned from vendor and needs to be installed.  Installed instrument and connected power,comm, and ground wires.  Verified instrument heaters are working and org_car = 4.88VDC on monitor.  Optical Scientific Inc.  Model 815DA, SN 05090146.  RTS at 1519 GMT.

9/27/11

PAD

UVSR building is needing outer skin replaced.  Facilities have to remove all power and communications boxes from outside of shelter to complete work.  They requested all instruments that power or comms run through UVSR shelter shut down.  At 1830 GMT, shut down Summinet and 2DVDisdrometer computers.  Removed all power and comm cables going to those instruments.  Network hub located in UVSR building was also shut down.  This affects the WRG and Disdrometer instruments.  RTS at 1930 GMT.

9/13/11

PAD

WRG bucket is full.  Bucket needs to be dumped and new antifreeze added.  Set flag at 1830 GMT.  Emptied bucket and added 1/2 gal of LowTox antifreeze and 1L of mineral oil.  Reset Tmm to zero.  RTS at 1850 GMT.

9/19/11

PAD

Disdrometer program show it is running correctly but processor did not pass daily test.  Comtrol serial/ethernet adapter located at processor does not have power.  Suspect lightening in recent weekend storms may have caused power supply to fail.  Troubleshooting found Comtrol adapter power supply has failed.  R/R power supply.  Processor will now communicate with computer.  RTS at 1940 GMT.

8/9/11

PAD

CF lost power on 8/8/11 at 2247 GMT.  Power was restored on 8/9/11 at 0211 GMT.  Found computer was inactive.  At 1519 GMT, set Disdrodata program to start measurements at 1600 GMT.   Able to connect with datalogger and transfer data.  RTS at 1534 GMT.

8/2/11

PAD

Mentor requests ORG instrument removed and sent for repairs and calibration.  Removed instrument at 1725 GMT.  Model ORG-815DA, SN 05090146.  RTS at 1745 GMT.

7/28/11

SD

Logger net window would not open.  Closed logger net and restarted program. RTS at 1635 GMT.

6/14/11

CAM

WBRG instrument vendor (Topper Nusbaum - Belfort) on site for maintenance.  Two of 3 transducers were failing.  Replaced all 3 transducers.  Replaced electronics and installed newest software version.  RTS status O.K. at 1430 GMT.

4/6/11

CAM

ORG and WRG collections are failing.  There is a window open stating Loggernet Server has encountered a problem and needs to close.  Rebooted computer and restarted program but same error occurred.  Disabled all collections in Loggernet Setup and the program would then stay running, but would lock up as soon as collections were enabled.  Under the Data Files tab in Setup, there is a table entry called DataSample and it was set for scheduled collections.  Disabled collections for this entry by unchecking the “Included for Scheduled Collection” box.  The following collections for WRG and ORG were successful.  RTS at 1320 GMT.

3/24/11

PAD

DQPR 2600 states WBRG sensor 2 is noisy.  Suggest that the connections to the transducers are checked.  Set port 1 flag at 1850 GMT and removed cover.  Reseated transducer inline connections and connections on control board.  No obvious problem found.  Removed flag and RTS instrument at 1905 GMT.

1/26/11

PAD

At 1642 GMT, unable to bring up connect screen for datalogger.  Closed out and tried to restart Loggernet, but error message shown:  Unable to start program because the Campbell Scientific communications server is already running.  Nothing is shown running in Task Manager.  Stopped Disdrodata program and rebooted computer.  Now able to connect with datalogger.  Set Disdrodata program to start measurements at 1700 GMT. 

12/7/10

PAD

At 1341 GMT, remotely connected with computer.  Computer acted sluggish and noticed “Setup” button on loggernet was depressed, but setup window was not opening.  Closed out loggernet and tried to restart but error was shown:  Unable to start program because the Campbell Scientific communications server is already running.  Rebooted computer and restarted disdrodata program.  RTS at 1401 GMT.

12/1/10

CAM

Site power needs to be shut off to install the transformer and run power to the new office building.  Instruments need to be shut down gracefully.  Stopped the programs and shut computer off at 1428 GMT.  Restarted after power was restored and RTS at 1845 GMT.  Set disdrometer measurement to start at 1900 GMT.

11/2/10

PAD

Last SDS collection was at 0501 GMT.  Unable to connect with ORG/WBRG datalogger.  Powercycled NL-100 interface.  Connected with datalogger and transferred data.  RTS at 1555 GMT.

11/1/10

PAD

Last SDS collection was at 0901GMT.  Unable to connect with ORG/WBRG datalogger.  Powercycled NL-100 interface at 1622 GMT.  Connected with datalogger and transferred data.  RTS at 1637 GMT.

10/28/10

PAD

Last SDS collection was at 0601 GMT.  Unable to connect with ORG/WBRG datalogger.  Powercycled NL-100 interface.  Now able to connect and transfer data.  RTS at 1451 GMT.

10/26/10

PAD

Unable to connect with ORG/WBRG datalogger.  At 1430 GMT, powercycled NL-100 interface.  Now able to connect and transfer data.  RTS at 1445 GMT.

10/20/10

PAD

Unable to connect with ORG/WBRG datalogger.  Powercycled NL-100 interface at 1337 GMT.  Now able to connect and data was transferred.  RTS at 1352 GMT.

10/14/10

PAD

Last SDS collection from ORG was at 0701 GMT.  Remotely accessed computer and was unable to connect with datalogger.  Rebooted computer at 1228 GMT with no change.  At 1431 GMT, powercycled NL-100 interface.  Able to now connect with datalogger and transfer data.  RTS at 1446 GMT.

10/1/2010

JEM

The disdrometer computer has been repaired and needs to be reinstalled.  Reconnected the computer and powered it on.  Made sure  the computer was able to communicate with the instruments and it was.  Was unsure of some of the collection settings, so notified the mentor and she logged in remotely to get everything properly set up.  RTS @1425 GMT.

9/29/10

PAD

At 1426 GMT, ORG computer shows error:  A problem has been detected and Windows has shut down to prevent further damage.  Powercycled computer and error shown on boot up:  Windows could not start because the following files are not found or corrupt:  \Windows\System32\Config\System.  Removed computer and computer dept suspects failed hard drive.  Troubleshooting at this time. 

9/27/10

PAD

Last SDS collection from ORG was at 0501 GMT.  Checked voltages at datalogger.  Powercycled NL-100 interface with no change.  Powercycled datalogger.  Able to connect with datalogger but connection failed after 1 min 8 sec.  No data was transferred.  Connected with datalogger using laptop and was able to connect and monitor ORG and WRG values.  Rebooted computer and was able to connect with datalogger but connection again failed after 1 min. 54 sec.  At 1512 GMT, found computer had power but nothing shown on monitor.  Powercycled computer and was able to connect with datalogger and transfer 15,975 values.  Set Disdrodata program to start measurements at 1600 GMT.  RTS at 1530 GMT.

9/10/10

CAW

TBRG removed from service and replaced with new WBRG.

5/13/10

PAD

CF lost power from 0623-0759 GMT.  At 1420 GMT, computer and instrument has power, but is not communicating.  Connected with datalogger and collected data. Set disdrodata program to start measurements at 1500 GMT.  RTS at 1435 GMT.

1/11/10

PAD

Last SDS collection was on 1/7/10 at 2205 GMT.  Unable to connect with disdrometer processor.  Troubleshooting with computer department found single port rocket port located at processor had failed.  R/R rocket port with one computer department had loaded firmware on and set up all communication variables.  Now able to connect with processor.  RTS at 2020 GMT. 

1/11/10

CAM

No data has been collected from the ORG or TBRG since 0800 GMT on 1/8.  Disdrodata program is unable to connect to CR1000 datalogger.  Used the status monitor in LoggerNet software to reset the devices PakBusPort_2 and LAN_2_CR1000.  Rebooted the computer and link between datalogger and computer was reestablished at 1700 GMT.  No data should be lost as all back data was being stored on the datalogger.

1/6/09

PAD

Last SDS collection for TBRG and ORG was 1020 GMT.  Unable to connect with datalogger.  Power cycled NL-100 interface, with no change.  When I returned to computer it shows:  A problem has been detected and windows has shut down to prevent damage to computer.  Multiple_IRP_Complete_Requests.  Restarted computer but still unable to connect with datalogger.  At 1745 GMT, was able to  remotely access computer and connect.  Suspect cold temperatures may be affecting some of the comms.  Troubleshooting at this time.  RTS at 1800 GMT.

1/6/10

PAD

Weekly tip test shows no tips.  T/S found rodents had cut signal wire from magnet to datalogger.  Repaired wire and repeated tip test.  RTS at 2204 GMT.

1/4/10

PAD

Disdrodata program has not been running since 12/31/09.  Error message was shown when trying to start program.  At 1543 GMT, started disdrodata program with no error shown.  Set program to start taking measurements at 1600 GMT.  Unknown at this time what was causing error message.  RTS at 1600 GMT.

12/31/09

PAD

At 1616 GMT, observed computer has power but nothing is shown on monitor.  Power cycled computer and monitor will now display.  Connected with datalogger for TBRG and ORG.  Error displayed when trying to start Disdrodata program:  There is a problem with Com 3.  (Error code 37).  Rebooted computer with no change in status.  Checked comms at processor with no issues found.  Power cycled processor with no change in status.  Notified instrument mentor via voice mail.  RTS at 1646 GMT.

12/10/09

PAD

SDS has not collected from TBRG since 0620 GMT.  Computer is unable to connect with datalogger.  At 1510 GMT, power cycled NL100 media convertor with no status change.  Power cycled data logger with no status change.  Rebooted computer and comms were established.  RTS at 1525 GMT.

12/10/09

PAD

SDS has not collected from ORG since 0701 GMT.  Computer is unable to connect with datalogger.  At 1510 GMT, power cycled NL100 media convertor with no status change.  Power cycled data logger with no status change.  Rebooted computer and comms were established.  RTS at 1525 GMT.

11/18/09

PAD

Connected ORG heater power wire to 12V on wiring panel as requested by mentor.  RTS at 1636 GMT.

10/19/09

PAD

Instrument returned from being repaired and mentor requests it placed in service.  Mounted ORG instrument, model ORG-815DA/ SN 05090146, and connected data/power and ground cables.  RTS at 1808 GMT.

9/22/09

MK

The ORG needs to be removed and sent back for factory repairs.  Removed ORG at 1535 GMT

9/15/09

PAD

Last successful collection was at 1601 GMT.  TBRG and ORG were also affected.  Unable to access Disdrometer computer remotely.  Able to ping server from Disdrometer computer.  Rebooted computer and was then able to remotely access computer.  RTS at 1815 GMT.

7/31/09

MK

Mentor has corrected CR1000 program but the data collections are unsuccessful. Discovered data directory names had defaulted and needed to be changed in order for collection script to have access to data from ORG. Changed all data collection folders to the original destinations. Site data system has resumed data collections from ORG at 1700 GMT.

7/29/09

MK

As per BCR 1577 the optical rain gauge needs to be rewired to operate in differential mode. This should clear up the false readings being reported as explained in DQPR 2173. Rewired ORG from 1915 to 1925 GMT. Loaded the new CR1000 program TipBucketOrgEventv3.cr1 onto data logger from ORG computer. Program load failed due to excessive characters on line 158. Mentor has been notified of problem.

7/15/09

PAD

DQPR 2173 requests check to make sure ORG instrument is properly grounded.  At 1556 GMT, removed all ground connections, cleaned and reseated.  No problems found.  RTS at 1620 GMT.

6/25/09

TMG

Computer software needs updated.  Stopped program @ 1928 GMT.  Installed Service Pack 3 @ 1929 GMT.   Rebooted and RTS @ 2011 GMT.

6/4/09

PAD

DQPR 2173 requests ORG heater power wire disconnected.  Suspect gnats are being attracted to heat source at night causing false readings.  Disconnected power wire to heater at 1838 GMT.  Notified mentor.  RTS at 1853 GMT.

4/2/09

MK

We are constructing a wind block around the Disdrometer in an attempt block the wind and verify that the intermittent readings are from high winds. Installed the cinder block enclosure at 1600 GMT. We removed it at 1516 GMT on 4/3/09.

3/30/09

PAD

SDS has not collected from TBRG since 3/27/09 at 1520 GMT.  At 1207 GMT, remotely accessed computer and forced collection from data logger.  Transfer progress showed 11 out of 12737 bytes transferred then it froze.  Disconnected/ reconnected with datalogger and tried collection again.  This time it showed 10437 bytes were transferred.  This would include data for the ORG as well.  RTS at 1222 GMT. 

3/30/09

PAD

SDS has not collected from ORG since 3/27/09 at 1601 GMT.  At 1207 GMT, remotely accessed computer and forced collection from data logger.  Transfer progress showed 11 out of 12737 bytes transferred then it froze.  Disconnected/ reconnected with datalogger and tried collection again.  This time it showed 10437 bytes were transferred.  This would include data for the TBRG as well.  RTS at 1222 GMT. 

1/13/09

PAD

Sensor is showing readings of small diameter droplets when it is not raining.  Troubleshooting has determined vibrations caused by wind is being detected by sensor.  Mentor requests something placed between sensor and base to isolate vibrations.  At 2114 GMT, stopped disdrodata program and installed pad between sensor and base.  Set disdrodata to start measurements at 2200 GMT.  RTS at 2129 GMT.

1/7/09

PAD

Disdrometer has been sporadically displaying rain droplets in the smaller diameter bins when it is not raining.  Mentor requests in DQPR 1980 instrument checked for contact between the splash guard support and the sensor housing where vibrations from splash guard caused by winds could be transferred to sensor.    At 1650 GMT, verified splash guard base is not contacting sensor cone guard.  Further troubleshooting found by lightly tapping on splash guard base, cone guard, or pallet base of instrument will produce false readings in smaller diameter bins.  Troubleshooting with mentor at this time about possible change in mounting of sensor.  RTS at 1705 GMT.

12/11/08

PAD

Raindrops are being displayed in smaller diameter bins on Disdrodata display when it is not raining.  Mentor requests cone on disdrometer plunger checked that cone has not loosed off of plunger threads causing movement and false readings. Checked cone for tightness at 1557 GMT and found no problems.  RTS at 1612 GMT.

12/9/08

PAD

Problem with Com 3 error is shown.  CF was without electricity between 1014 and 1222 GMT.  Started disdodata program to start measurements at 1600 GMT.  RTS at 1604 GMT.

7/14/08

MK

CF lost power on 7/13/08 between 0028 and 0137 GMT. Restarted CR1000 and Disdrometer software at 1511 GMT.

7/11/08

PAD

TBRG funnel is obstructed with insects.  Water is being held back in funnel from recent rains.  Cleared obstruction and emptied funnel.  RTS at 1510 GMT.

7/3/08

PAD

Troubleshooting droplets being measured up to .4 mm when it is not raining concerning DQPR 1980.  Cone was removed on 7/1/08 and needs to be installed on sensor.  Installed sensor cone and surrounding guard at 1433 GMT.  Started measurements at 1500 GMT.

7/1/08

MK

DQPR 1980 indicates instrument is registering raindrops up to 0.4 mm in size when no rain is present.  Suspect environmental factors are causing noise in data.  Checked instrument for obvious causes of noise and found none.  Removed cone and covered instrument to verify noise in data is being caused by electronics or environment, wind and insects.  RTS at 1600 GMT.

4/14/08

PAD

SDS has not collected since 4/11/08 at 2002 GMT.  Remotely access computer at 1353 GMT.  Unchecked “Pause all scheduled tasks” box.  Computer will now collect data from datalogger.  RTS at 1408 GMT.

4/8/08

PAD

Power outage at CF between 0709 and 0910 GMT.  Computer has power but nothing is shown on monitor.  Unable to get anything on monitor.  Tim Grove removed computer and put hard drive into another computer.   At 1842 after installing computer, reset media converters and restarted disdrodata program.  TBRG and ORG were also affected during this time.  RTS at 1857 GMT.

3/10/08

CAM

The ORG has been falsely indicating precipitation when the temperature is below freezing and the humidity is high.  The heaters do not appear to be working, as the lenses do not feel warm.  Contacted the vendor who suggested returning the ORG for warranty repair.  Powered the ORG off at 1535 GMT.  R/R *ORG.  Powered new ORG on and RTS at 1547 GMT.  Felt the lenses after a few minutes and they felt slightly warmer than ambient, but it is a very subtle difference. 

*Old – 07060272  *New - 05090146

3/6/08

CAM

Instrument mentor requested the ORG heater power line be checked.  Verified that the heater power wire was connected to 12V on the datalogger.  Also disconnected the cable from the ORG at 2105 GMT and measured 13VDC on pin J of the cable.

2/15/08

DN

Check the ORG alignment.  The path between the transmitter and the receiver was found to lie on a bearing of 15 degrees magnetic, which is about 10 degrees east of true north for the SGP CF.  I corrected the alignment so that the optical path was oriented along a true N/S line.  The sensor was level.  RTS 1655 GMT.

12/5/07 – 12/6/07

CAM/PD

DQPR 1868 indicates the ORG has been reporting constant precipitation of around 0.1 mm/hr.  Checked the output of the ORG at the datalogger and found it to be around 87mV, which would result in the 0.1 mm/hr reading.  The output of the ORG should be less than 85mV during non-precipitation.  Checked all grounding connections and found the datalogger earth ground had to be grounded at the same point as the ORG sensor’s earth ground in order to read correctly.  Ran a temporary ground wire from the datalogger earth ground to the ORG sensor’s ground rod and the output dropped to a steady 41mV, which results in 0 mm/hr precipitation reading.  Facilities personnel will acquire a suitable ground wire for permanent burial and complete the install later.  For now, the ground wire is run above ground and the instrument is operating correctly.  RTS at 2100 GMT.

10/16/07

DN

ORG installed and connected to the RAIN CDL.

10/3/07

CAM

The site lost power for an extended time beginning around 1940 GMT on 10/2 during a severe thunderstorm. Instrument ran for a short time on UPS power and then shut down.  Power was restored after normal business hours yesterday.  Computer was operational this morning but indicated an error with COM 3.  Closed the error message and started Disdrodata and LoggerNet at 1403 GMT.  RTS status O.K.

7/18/07

PAD

Last SDS collection for the TBRG was on 7/17/07 at 1204 GMT.  Computer shut down due to loss of UPS.  Mentor remotely reset IP address for TBRG datalogger CR1000 at 1535 GMT.  Data file is not being created for TBRGeng.dat.  At 1705 GMT, notified instrument mentor.  Mentor is troubleshooting at this time. 

7/17/07

PAD

Computer has lost connection with disdrometer processor and TBRG datalogger.  Reset media converters at instrument and computer.  Computer will not connect with datalogger.  Reset media converters again, when I returned to computer, error message was displayed:  Multiple_IRP_Complete_Requests.  It said computer had encountered problems and had shut down Windows.  Could be caused by recent software download.  If this is the first time to see error, reboot computer.  Rebooted computer and reset media converters again.  Started disdrodata collection to start at 1900 GMT.  Computer will not connect with TBRG datalogger.  Contacted instrument mentor.  Troubleshooting at this time.  RTS at 1930 GMT.

5/29/07

MK

The power at the CF had failed some time over the weekend causing the UPS to shutdown due to low voltage on batteries. Powered on UPS at 1434 GMT, restarted Distrometer computer and software at 1440 GMT.

4/30/07

CAM/PAD

The disdrometer came back from the vendor and needed to be reinstalled.  Installed instrument and processor at 1645 GMT.   Set program to start at 1700 GMT.  RTS status O.K.

3/7/07

CAW

Removed rain gauge from service for calibration at 1430 GMT.  Replaced same gain gauge and RTS at 2000 GMT.  

2/26/07

PAD

Instrument removed from service and returned to vendor for calibration.

8/28/06

PAD

Last successful collection was on 8/25/06 at 2106 GMT.  On 8/28/06, found UPS that supplies power to instrument had tripped.  Reset UPS and media converters between instrument and computer.  UPS had not tripped until sometime on 8/26/06. At 1402 GMT, restarted computer and set program to start measurements at 1500 GMT.  RTS at 1417 GMT.

5/9/06

TMG

Stopped the program at 2010 GMT.  Edited registry settings to give SDS full control over net time.  Rebooted computer and RTS at 2023 GMT.

4/14/06

Mentor

System installed by mentor.